# Inner crust of neutron stars: Unified description of structure and superconductivity

### v1

@@ I Introduction
The physics of neutron star interiors represents a major challenge for both nuclear physics and astrophysics. There are robust theoretical models for describing the structure of neutron stars up to densities of the order of 1-2 times the the saturation baryon number density in symmetric nuclear matter ($n_{0}\sim 0.16$ fm-3). However, in the central part of the core of neutron stars the matter is expected to be at significantly higher densities. At present, no reliable theoretical model exists to describe matter at such high density. Fortunately, observations of neutron stars suggest physical phenomena that likely probe the global structure of neutron stars including the central part of the core. Among those are the giant magnetar flares and the corresponding quasi-periodic oscillations in the afterglows. As is apparent from the energetic scale of the phenomenon, it involves hydromagnetic waves spreading across the entire star.
Understanding the hydromagnetic waves on a global stellar scale requires to specify properties of the electrical conduction of matter and its interaction with the magnetic field. In this paper, we will study the nature of the electrical conductivity in the inner crust with emphasis on superconductivity of protons and the microscopic structure of the pasta phases.
The pasta phases appear in the bottom part of the inner crust as nonspherical nuclear clusters CaplanHorowitz2017 . Since the proton pairing is generally expected in nuclear matter around the nuclear saturation density BPP1969 , the protons inside these nonsperical structures should be superconducting Kobyakov2018 ; KobyakovPethick2018 ; ZhangPethick2021 . However, it is now well-known LimHolt2021 that magnitude of the proton pairing gap is very sensitive to both the baryon density and the proton fraction. Thus, understanding of the essential details of superconductivity in the pasta phases requires a complete microscopic description of the structure of the inner crust of neutron stars.
Information regarding the microscopic structure of the pasta phases is crucial for the correct choice of the superconductivity model, which is theoretically applicable for the pure structures of the pasta phases. For instance, let us consider the 1-dimensional pure pasta structure (lasagna phase, or 1N phase), where the slabs of superconducting nuclear matter are separated by regions containing the pure neutron liquid. The distance between the nearest interfaces of two adjacent slabs is denoted by $d_{L}$, while the width of the slab is $2r_{N}$. Under the expected physical conditions, such lasagna phase is similar to a layered superconductor and should be described by different mean-field models depending on the microscopic parameters. When the tunneling of Cooper pairs between the slabs through the insulating pure neutron liquid is small, the appropriate description is given by a discreet (in the sense that the slabs are individual) model of superconductivity, for example, the Lawrence and Doniach model. On the contrary, when the tunneling between the slabs is significant, the system should be described as a continuous (in the sense that the superconducting density is averaged across many slabs) model of superconductivity, for example, the Ginzburg-Landau (GL) model KlemmEtal1975 ; DeutscherEntinWohlman1978 ; WangEtal2001 .
As is well-known KlemmEtal1975 ; DeutscherEntinWohlman1978 ; WangEtal2001 , the significance of the Cooper pair tunneling is quantified by the relative magnitude of the coherence length $\xi$ associated with the uniform matter inside the superconducting slab and the separation distance $d_{L}$. A continuous model for superconductivity can be used when $\xi\gtrsim d_{L}$, while at $\xi\lesssim d_{L}$, the discreet model should be preferred.
The continuous model is generally valid at temperatures close to the superconducting critical temperature (in this case $\xi\rightarrow+\infty$), but may be invalid in the low-temperature regime $T\rightarrow 0$ well below the critical temperature, if $\xi<d_{L}$. In our system though, the open question at present is whether the superconducting critical temperature is smaller or larger than the lattice melting temperature. In a recent work, Zhang and Pethick ZhangPethick2021 have used a continuous model for superconductivity in the lasagna phase and implicitly assumed that the superconducting gap energy is constant across the entire lasagna phase.
We are interested in a quantitative description of the energy landscape for the equilibrium pasta structures taking into account superconductivity of protons. While the magnetic properties of cold neutron stars have long been recognized as an observational clue to the internal stellar structure, however, their understanding has been hindered by the possible presence of superconductivity near the crust-core boundary. Another strong source of complication is the possible presence of the pasta phases, because those phases are anisotropic and multiply connected locally, in contrast to the isotropic locally (apart from the global stellar length scale) singly-connected superconductor in the outer core, which makes their magnetic properties very different from those of the “simple” bulk superconductor in the outer core.
A comprehensive study of the inner crust including the pairing correlations should be performed on the basis of the complete thermodynamic potential (the Gibbs free energy) including such variables as temperature, entropy per particle, the magnetic induction, the given (external) magnetic field and the baryon number density. In this path, we will make the first step and study the internal energy per baryon in the pasta phases. Also we will establish a basis for further studies of superconductivity by studying of the superconducting coherence length and the pairing gap of protons in the pasta phases. Effects of entropy and the magnetic field will be studied in the future work.
@@
Within a mean-field approach, the most sophisticated description of the inner crust is provided by Hartree-Fock (HF) calculations in the Wigner-Seitz (WS) cells, which are complicated owing to the presence of the neutron liquid and the need to deal with non-spherical geometries. There are many Hartree-Fock calculations of spherical nuclei since the seminal paper of Negele and Vautherin negele73 (see also mondal20 and references therein) and also more recently 3-dimensional mean-field calculations using non-relativistic and relativistic energy density functionals and classical and quantum molecular dynamics simulations (see for further details the recent paper by Newton et al. NewtonEtal2022 where the state-of-the art of nuclear pasta calculations is discussed in detail). However, these sophisticated calculations only provide partial information about the structure of the inner crust and to date no systematic calculations to obtain the complete equation of state in the inner crust have been made. Large scale calculations of the inner crust are usually performed with the the help of semiclassical approaches based on the Thomas-Fermi (TF) approximation (see sharma15 for more details) or through the LDM.
In this paper we will work within compressible liquid drop model (LDM), which is a powerful tool offering a simple physical interpretation of complicated many-body nucleon effects. This model is based on a convenient choice of the collective degrees of freedom. It has been used in the pioneering study of the nuclei surrounded by dripped neutrons BBP1971 and has been crucial in the theoretical prediction of the nonsperical nuclei in the inner crust of neutron stars RavenhallEtAl1983 ; HashimotoEtAl1984 .
@@
Within the LDM, the properties of the nuclear surface, as for instance the surface tension and the behavior of the neutron and proton density profiles, play an essential role. The surface tension has been always represented in approximate manner, based, for example, on the leptodermous expansion of the binding energy of nuclear systems myers69 ; myers74 . The orders of approximation correspond to the power of the moment in the integrand in Eqs. (191), (193) and (195). The leading approximation is to calculate the surface tension associated with the planar surface (when the principal curvature is zero) with a sharp density profile of protons and use it for description of the nuclear clusters. The approximate character of the surface tension in this case comes from the fact that in reality, the principal curvature of the surface of nuclear cluster is nonzero, the proton surface is smeared, and there is a film of adsorbed neutrons at the surface of nuclear cluster. In the next-to-leading approximation, the energy contribution due to a nonzero principal curvature is included, while the proton surface is kept sharp and no neutrons adsorbed; this approximation is elaborated in this paper. In the next orders of approximation, which are left for the future work, the smearing of the proton surface and neutrons adsorbed on the nucleus surface are included.
The next-to-leading approximation was considered in many recent LDM calculations newton13 ; LimHolt2017 ; carreau20a ; carreau20b ; dinh21a ; dinh21b ; balliet21 ; shchechilin22 . Importantly, in those calculations the surface and curvature contributions to the energy density functional were taken from parameterizations originally proposed by Lattimer, Pethick, Ravenhall and Lamb ravenhall83 ; lattimer85 generalized later on by Lorentz and Pethick lorentz93 . Here, we improve on the earlier calculations and extract the planar and curvature contributions to the surface tension from a self-consistent extended Thomas-Fermi (ETF) brack85 calculation performed using the same interaction that describes the bulk part in the LDM, following closely the method described in detail in Refs. Vinas1998 ; douchin00 , which is summarized in Appendix D.
@@
The potential energy contribution to the energy density functional provided by Skyrme forces can be expressed in terms of the nucleon densities in uniform nuclear matter $n_{n}$ and $n_{p}$, and the square of their gradients, $({\bf\nabla}n_{n})^{2}$ and $({\bf\nabla}n_{p})^{2}$. In the ETF approach, the kinetic energy is also an energy density functional that is expanded in powers of 2superscriptPlanck-constant-over-$\hbar^{2}$. In this work we take only the two first terms of this expansion. The leading order term correspond to the bulk, which is expressed through the TF kinetic energy densities for neutrons and protons. The next-to-leading order term includes second order gradients of the nuclear densities and effective masses as well as a contribution of the semiclassical spin density, as it can be seen in Eqs. (35), (36) and (37).
The paper is organized as follows. Section II presents basic equations of the LDM of nuclear matter, which are explained in detail in Appendices B and C. Appendix A lists the details of the nuclear interaction energy functional used in the paper and a discussion of the nuclear matter properties. Appendix D describes our self-consistent numerical results based on a single equation of state for both the nuclear bulk and the surface properties. Section III is reserved for the numerical results on solution of the LDM equations, evaluation of thermal effects, the workfunction to transfer proton from nucleus to the pure neutron matter, the coherence length of superconducting protons and the magnitude of the pairing energy gap. In Section IV we analyze the superconductivity in the 1-dimensional pasta phase by studying the coherence length of superconducting protons and the stress tensor in the superconducting-superfluid mixture relevant to the outer core and the inner crust of neutron stars. Section V summarizes the paper.
II Liquid drop model
II.1 Basic approximations
We assume that the matter in the inner crust of neutron stars is in a low-energy static state and the nuclei form a Coulomb crystal with a uniform electron background. Therefore, it is a reasonable approximation to represent the real matter as a model with ideal lattice and nuclei with sharp edges. Inside these nuclei, the matter is assumed to have a uniform density of baryons $n$ with a local number density of protons $n_{p}=xn$, $x$ being the proton fraction, and the local density of neutrons $n_{ni}=(1-x)n$. The voids between the nuclei are filled by the dripped neutrons with a uniform density $n_{no}$.
@@ II.1 Basic approximations
We shall work in the spherical Wigner-Seitz (WS) cell approximation and focus on a single unit cell with volume $V_{c}$ and the total number of neutrons $N_{n}$ and protons $N_{p}$:
@@
A cell either contains a nucleus surrounded by the neutron liquid (1N, 2N, 3N), or it is entirely filled by the nuclear matter except in its center, where there is a bubble filled with a pure neutron liquid (1B, 2B, 3B). The nuclear matter and the pure neutron liquid are separated by a spherical interface in case of 3N and 3B, by a cylindrical interface in case of 2N and 2B, and by a planar interface in case of 1N. The WS cells have a finite lattice period along three spatial dimensions (for 3N and 3B configurations), or along two dimensions (for 2N and 2B configurations), or along only a single dimension (for 1N configuration). The length or the area of the cell along the direction with a continuous translational symmetry is equal to the unit of length or the unit of area, respectively.
@@
Notice that in the one-dimensional symmetry, the description in terms of 1N is equivalent to the description in terms of the 1B with an appropriate change in variables, and thus, the results for 1N and 1B can be combined and represented as a single case, which is denoted briefly as 1N. Still, it is important that the volume fraction for 1N corresponds to the ratio between the volume of nucleus and the unit cell, while the volume fraction for 1B corresponds to the ratio between the volume of neutron-filled bubble and the unit cell, hence the variable $u$ used for the volume fraction in the phase of nuclear clusters is distinct from the variable $u^{\rm bub}$ used for the volume fraction in the phase of nuclear bubbles. To keep the theory as clear as possible, in Appendix we present the formulas for nuclei and for bubbles separately.
II.2 Basic equations

### v2

# Inner crust of neutron stars: Polymorphism and superconductivity in the liquid drop model

@@ I Introduction
Description of the problem. The physics of neutron star interiors represents a major challenge for both nuclear physics and astrophysics. The ultimate question is: What is the structure of matter inside neutron star? There are robust theoretical models for describing the structure of neutron stars up to densities of the order of 1-2 times the saturation baryon number density $n_{0}$ in symmetric nuclear matter, which is of the order of $0.16$ fm-3. However, in the central part of the core of neutron stars the matter is expected to be at significantly higher densities. At present, no reliable theoretical model exists to describe matter at such high density. Fortunately, observations of neutron stars suggest physical phenomena that likely probe the global structure of neutron stars including the central part of the core. Among those are type-II supernova events, which involve a collapsing stellar core with the relevant matter energy density BrownBetheBaym1982 and the gravitational wave observations OyamatsuIidaSotani2020 . Another example is provided by giant magnetar flares and the corresponding quasi-periodic oscillations in the afterglows observed in X-ray astronomy TurollaZaneWatts2015 . It is apparent from the energetic scale of these phenomena and from the corresponding basic physical picture of neutron star observable dynamics, that all these phenomena involve some kind of hydromagnetic waves spreading across the entire star. Understanding the physics of hydromagnetic waves on a global stellar scale requires, first of all, to specify structure of the neutron star matter and properties of the electrical conduction of matter that determines its interaction with the magnetic field.
Since the pioneering quantitative studies of structure of the neutron star matter carried out by Baym, Bethe and Pethick BBP1971 , much work has been done in this direction. It is convenient to categorize the studies according to the relevant length scale. Here, we denote by the term microscopic structure the nuclear matter structure at scales between $\sim 1$ fm and $\sim 50$ fm. In contrast, the term macroscopic structure is understood in this work as the length scales larger than $\sim 50$ fm, corresponding to thermodynamic subsystems which include many unit cells of the crystallized matter of the crust, assuming that for such subsystems the thermodynamic description (that is, for instance, involving temperature) is appropriate.
A notable step in microscopic studies was done in RavenhallEtAl1983 ; HashimotoEtAl1984 , unveiling the theoretical prediction that nuclear matter in the neutron star crust might exist in the form of nonsperical nuclei immersed in the pure neutron liquid. Geometrical structure of such matter resembles the well-known varieties of ordinary pasta such as lasagna and spaghetti. As a result, the term pasta phases has been adopted in nuclear astrophysics for the nonsperical nuclei in neutron star CaplanHorowitz2017 . In the present work, we will focus on the inner crust matter, which is expected to host pasta phases.
Structure of the paper. In Section II we present more details on the problem considered in this paper and provide a brief review of the relevant literature. Liquid drop model used for description of the nuclear properties is explained in Section III with additional details presented in Appendices A-D. In Section IV we report the numerical solution to the equations of liquid drop model. In Section V we propose a simple model for superconductivity incorporated into liquid drop model. Using our numerical solution, in Section VI we analyze effects of temperature on the expected structure of the inner crust and evaluate some characteristic quantities of superconducting lasagna. Section VII is devoted to study of stress in a setting where a spherical layer of lasagna is subject to a pulsar spin lag between the charged particles and superfluid neutrons.
II Overview of studies of structure of pasta phases
Studies of microscopic structure. Further studies of the microscopic structure of pasta phases have revealed that the energy differences per baryon between the different phases are quite small LattimerEtal1985 . It has been found that the quantum corrections to the ground-state energy in neutron matter (the shell effects associated with unbound neutrons) are important and provide considerable uncertainty for understanding of the realistic structure of the crust BulgacMagierski2001 ; MagierskiHeenen2002 . The effect originates both from structure at the microscopic scales where the precise geometry of Wigner-Seitz cell is defined, and from the macroscopic structure that contains information about the spatial distribution of structural lattice defects.
@@ II Overview of studies of structure of pasta phases
Within the mean-field quantum description based on the Schrodinger equation, the most technically sophisticated approach is provided by Hartree-Fock calculations (or Hartree-Fock-Bogoliubov equations if pairing is included). There are many Hartree-Fock calculations of spherical nuclei since the seminal paper of Negele and Vautherin negele73 (see GrillMargueronSandulescu2011 ; ShelleyPastore2020 and also mondal20 and references therein). Lately, 3-dimensional mean-field calculations using non-relativistic and relativistic energy density functionals and classical and quantum molecular dynamics simulations have been performed NewtonEtal2022 . Also, 1-dimensional time-dependent density functional theory accounting for band effects and superfluidity of neutrons has been established SekizawaEtal2022 ; YoshimuraSekizawa2024 However, those calculations only provide partial information about the structure of the inner crust, because the problem of finding pasta phases with the minimum energy in the entire range of pressures corresponding to the inner crust has not been addressed in the full quantum approach. Recently, the quantum effects have been studied in the spirit of Strutinsky energy theorem, as a perturbative correction to the quasiclassical Thomas-Fermi calculation in mondal20 and PearsonChamel2022 . Thus, to date, no full quantum description to obtain the complete equation of state in the inner crust has been performed.
In studies of the structure, it is reasonable to start from liquid drop model BBP1971 , which allows to clearly separate different contributions to thermodynamic quantities and paves the way to understanding various phenomena physically LattimerEtal1985 . In liquid drop model, theoretical uncertainties stemming from the quasiclassical nature of the model and from the uncertainty of the input nuclear energy functionals can be relatively easily evaluated and improved in a systematic way.
Microscopic description of the inner crust structure is also performed with the help of semiclassical approaches based on the Thomas-Fermi approximation (see sharma15 for more details) or through liquid drop model. The latter does not involve the Schrodinger equation, but nevertheless is capable of a microscopic description on length scales of a single nucleus. In addition, liquid drop model can be applied for inquiry regarding macroscopic, thermodynamic effects.
In this model, the properties of the nuclear surface, as for instance the surface tension and the behavior of the neutron and proton density mean-field profiles, play an essential role. The surface tension has been always represented in approximate manner, based, for example, on the leptodermous expansion of the binding energy of nuclear systems myers69 ; myers74 .
@@
In the leptodermous expansion, the orders of approximation correspond to the power of the moment in the integrand in Eqs. (175) and (177). The leading approximation of liquid drop model is to calculate the surface tension associated with the planar surface (when the principal curvature is zero) with a sharp density profile of protons and use it for description of the nuclear clusters. The approximate character of the surface tension in this case comes from the fact that in reality, the principal curvature of the surface of nuclear cluster is nonzero, the proton surface is smeared, and there is a thin surface layer of adsorbed neutrons at the surface of nuclear cluster. In the next-to-leading approximation, the energy contribution due to a nonzero principal curvature, the proton surface smearing and the neutrons adsorbed at the surface are included. In this paper, we have only considered the curvature term, neglecting the two other contributions, whose impact is rather small.
Next-to-leading approximations of liquid drop model were considered in many recent liquid drop model calculations NakazatoIidaOyamatsu2011 ; newton13 ; LimHolt2017 ; carreau20a ; carreau20b ; dinh21a ; dinh21b ; balliet21 ; shchechilin22 . Parametrization of the planar surface tension in liquid drop model is often borrowed from the results of Lattimer, Pethick, Ravenhall and Lamb ravenhall83 ; LattimerEtal1985 , which were generalized later on by Lorentz and Pethick lorentz93 . Interestingly, in many earlier calculations the curvature was neglected, while the planar surface contribution and the bulk contribution were taken from different parametrizations of the nuclear energy, raising questions about self-consistency.
@@
In this paper, we improve on the earlier calculations and extract the planar and curvature contributions to the surface tension from a self-consistent extended Thomas-Fermi brack85 calculation performed using the same interaction that describes the bulk part in liquid drop model. To this end we follow closely the method described in detail in Refs. Vinas1998 ; douchin00 , which is summarized in Appendix D. The potential energy contribution to the energy density functional provided by Skyrme forces can be expressed in terms of the nucleon densities in uniform nuclear matter $n_{n}$ and $n_{p}$, and the square of their gradients, $({\bf\nabla}n_{n})^{2}$ and $({\bf\nabla}n_{p})^{2}$, which simulate the finite range of the nucleon-nucleon interaction. In the extended Thomas-Fermi approach, the kinetic energy is also an energy density functional that is expanded in powers of $\hbar^{2}$. In this work, we take only the two first terms of this expansion. The leading order term corresponds to the bulk, which includes the Thomas-Fermi kinetic energy densities for neutrons and protons. The next-to-leading order term contains the second order gradients coming from the Skyrme force and the Weizsacker correction to the kinetic energy density and effective masses, as well as a contribution of the semiclassical spin density, as it can be seen in Eqs. (47), (48) and (49). We now turn to studies of the macroscopic structure.
Studies of macroscopic structure. The theoretical development has led to realization that the actual state of the crust is unlikely to be a system of stratified layers of pure pasta phases, but rather is likely to be a collection of monocrystals Jones1990 ; MagierskiHeenen2002 ; KobyakovPethick2015 . The notion of polycrystalline matter that is constructed from monocrystals separated by planar structural defects, sounds plausible and is usually considered quite obvious Jones1990 by virtue of the analogy of the neutron star solid crust with known geological solids. Still, it is expected that the mechanical nature of the pasta phases is much more complex than that of polycrystalline matter. In fact, only the phases with spherical nuclei (or spherical bubbles, if exist) solidify into a solid state, while other pasta phases such as the spaghetti, lasagna and their bubble counterparts are very similar to liquid crystals PethickPotekhin1998 ; KobyakovPethick2018 . Moreover, since the proton pairing is generally expected in nuclear matter near the nuclear saturation density BPP1969 , the protons inside the nonspherical nuclear clusters of the pasta phase should be superconducting Kobyakov2018 ; KobyakovPethick2018 ; ZhangPethick2021 . Thus, the most striking physical difference with the usual liquid crystals is, perhaps, superconductivity of protons in the dense clusters of the pasta phases.
However, establishing further details in physical picture of the macroscopic structure is a complex task CaplanHorowitz2017 . There are many factors which generate uncertainty within description of the crust structure on macroscopic scale. For instance, it turns out that the stellar magnetic induction as moderate as $\sim 10^{15}$ G is important in calculations of the crust structure as soon as the proton pairing is taken into account Kobyakov2024 . In this case, the proton pairing in the magnetic field leads to generation of the Meissner electric currents due to superconducting proton flow. The latter tend to screen the magnetic field and to bunch its lines into the quantized magnetic vortices of proton fluid, while the electron magnetism can be neglected. In lasagna, the ability of the material to develop the Meissner proton currents is clearly dependent on the relative angle between spatial orientation of lasagna structure and the direction of the magnetic field. Some characteristic cases are demonstrated in panels (a) and (b) of Fig. 8 below. The reason is that the superconducting electric currents might be impeded by the pure neutron matter, that fills the space between the nuclear clusters. Therefore, energy of the interaction between the magnetic field and lasagna has an angular dependence. This implies that there is an angular variation of the interaction energy. The initial calculations in Kobyakov2024 have shown that this variation of the interaction energy is of the same order as the energy differences between various pasta phases, thereby pointing to an important role of a moderate strength magnetic field of $\sim 10^{15}$ G for investigation of the realistic structure of the inner crust.
Information regarding the superconducting gap energy and the microscopic structure of the pasta phases is crucial for the correct choice of the superconductivity model. The latter is necessary for description of the magnetic properties of the pasta phases and, thus, is relevant for the astrophysical applications of the theory. In lasagna, different superconductivity models correspond to qualitatively different physical situations, depending on the relation between strength of the proton pairing energy for two protons both from the same nuclear cluster or for two protons from the neighboring clusters. From this viewpoint, the system is analogous to the terrestrial high-temperature superconductors. The latter is a class of superconducting materials physically structured as a stack of thin electrically conducting layers separated by thin insulating layers KettersonSong1999 . From experimental studies of the high-temperature superconductors, it is known that in case when a coherence length of the superconducting order parameter associated with quasi uniform matter inside the conducting layers is smaller than the distance between the edges of the conducting layers, then a discrete model of superconductivity is required. A standard tool in this case is the Lawrence and Doniach model (see Clem1992 and references therein). In the opposite case when the coherence length is larger than the distance between the edges of the conducting layers, the physical situation is qualitatively different and the appropriate description is given by the anisotropic Ginzburg-Landau (or, continuous) model KlemmEtal1975 ; DeutscherEntinWohlman1978 ; WangEtal2001 .
As a matter of fact, the continuous superconductivity model is generally valid at temperatures close to the superconducting critical temperature, if the superconducting transition temperature $T_{c}$ is smaller than the pasta crystallization temperature $T_{m}$. This behavior is expected just below $T_{c}$, because in this case the coherence length $\xi$ becomes very large and easily exceeds the distance $d_{L}$ between the clusters. However, the continuous model might be invalid in the low-temperature regime well below the superconducting transition temperature, if $\xi$ becomes smaller than $d_{L}$.
Finally, it is worth to note that the discreet superconductivity model was assumed in Kobyakov2018 . One of the most interesting consequences for such assumption is that, in the absence of bridges or proton tunneling connecting the neighboring slabs in lasagna, the penetration depth of the magnetic field which is parallel to the slab surface, is infinite. This, in turn, implies that under some symmetric conditions (see Fig. 10 below), the magnetic stress might well exceed the yielding stress of the crust material Kobyakov2018 . In the further work, Zhang and Pethick ZhangPethick2021 have used a continuous model for superconductivity in lasagna phase.
In this work, we will introduce a simple model of the superconductivity in lasagna and extract the density-dependent value of superconducting energy gap inside the nuclear cluster of lasagna from the calculations of LimHolt2021 for uniform nuclear matter. As a result, we shall demonstrate that in defect-free lasagna, a crossover between the discreet and the continuous superconductivity regimes might take place at some relevant pressure. Certainly, such approach is only a rather crude approximation, because (i) in lasagna, the superconducting gap might be anisotropic in the momentum space even if S-wave pairing for neutrons is assumed and (ii) matter inside the clusters of lasagna is not homogeneous EstalEtal2001 ; PankratovEtal2008 ; SchuckVinas2011 ; OkamotoEtal2013 ; FurtadoEtal2022 ; YoshimuraSekizawa2024 . However, we think that this simplified approach is a reasonable starting point for understanding of physics for superconductivity effects in magnetic field.
III Description of liquid drop model
III.1 Basic approximations
In order to study quantitatively the microscopic structure of nuclear matter in the conditions of the inner crust of neutron stars, we first consider a theoretical model, in which the matter is assumed to be in its ground state at zero temperature. This implies an ideal crystal lattice, which is conveniently described in terms of a minimal unit cell (or, Wigner-Seitz cell, or simply the unit cell). We also assume that the electrons are uniform at the scale of a single unit cell, which is a good approximation to the realistic state of the electrons due to the fact that their many-body state is degenerate and ultrarelativistic. The baryon density averaged at length scales much larger than size of the unit cell is denoted as $n_{b}$. Locally inside the nuclear cluster, the matter is assumed to have a uniform density $n$ associated with baryons with a local number density of protons $n_{p}=xn$, $x$ being the proton fraction, and the local density of neutrons $n_{ni}=(1-x)n$. The voids between the nuclei are filled with the dripped neutrons with a uniform density $n_{no}$.
@@ III.1 Basic approximations
We focus on the unit cell with volume $V_{c}$, which contains the total number of neutrons $N_{n}$ and protons $N_{p}$:
@@
A cell either contains a nucleus surrounded by the neutron liquid. In case of the bubble phases, the unit cell is entirely filled with the nuclear matter except in its center, where there is a bubble filled with a pure neutron liquid. Correspondingly, the types of unit cell are called pasta phases and are denoted as 1N (lasagna), 2N (spaghetti), 3N or, for the bubble phases, 1B (lasagna), 2B, 3B.
@@
In case of 3N and 3B, the nuclear matter and the pure neutron liquid are separated by a spherical interface. Correspondingly, in case of 2N and 2B the interface is a cylindrical surface with infinite length. In 1N and 1B phases, the interface is planar surface with infinite area. The unit cells have a finite lattice period along three spatial dimensions (for 3N and 3B), or along two dimensions (for 2N and 2B), or along only a single dimension (for 1N and 1B).
@@
Notice that in 1N and 1B phases, the picture in terms of 1N is equivalent to the picture in terms of 1B with an appropriate change in variables, and thus, the results for 1N and 1B are combined and represented as a single case, which is denoted briefly as 1N (or, lasagna). Still, it is important that the volume fraction for 1N corresponds to the ratio between the volume of the nucleus and the unit cell, while the volume fraction for 1B corresponds to the ratio between the volume of neutron-filled bubble and the unit cell. Hence the variable $u$ used for the volume fraction in the phase of nuclear clusters is distinct from the variable $u^{\rm bub}$ used for the volume fraction in the phase of nuclear bubbles. To keep the theory as clear as possible, in Appendices B-C we present the formulas for nuclei and for bubbles separately. Finally, we note that we use the most symmetric approximation for the unit cell (such as the spherical unit cell in 3N and 3B phases), which leads to certain relation between the volume fraction and the nucleus radius, see Eqs. (63) and (123).
III.2 Basic equations and scheme of numerical solution